OpenStudy (anonymous):

the length of a rectangle is four times its width .the area of the rectangle is 100 square centimeters.find the dimensions of the rectangle.

OpenStudy (ddcamp):

A = l * w l = 4w A = 4w*w → 4w² 100 = 4w² 25 = w² 5 = w 100 = 5 * l 20 = l

OpenStudy (anonymous):

thanks